OpenStudy (anonymous):

How do you create a formula to calculate the number of people in the U.S. if 1/4th of the people equals 57.7 million? and it has to be done on excel

OpenStudy (anonymous):

\[x/4=57.7million\]

OpenStudy (anonymous):

it has to be done on ecel and i have no clue how to do that either nor gragh that

OpenStudy (anonymous):

sorry *graph, *excel

OpenStudy (anonymous):

there is nothing to graph. \(\frac{x}{4}=57.7\) means \(x\) is a number, namely \(4\times 57.7\)

OpenStudy (anonymous):

i under stand that much but it has to be done on excel and they are asking for a graph based on the materail of "working statistics" which is saying how much people rather be doing other things than atending counciling and in the finding it has percentiles